OpenStudy (anonymous):

The formula for the volume of a sphere is V=4/3 pi r cubed. What is the radius of the basketball?

OpenStudy (turingtest):

you are given the volume of the ball I assume?

OpenStudy (anonymous):

No you are not

OpenStudy (turingtest):

well then you can only get the answer in terms of V you are given that\[V=\frac43\pi r^3\]can you solve this for \(r\) ?

OpenStudy (anonymous):

I believe i did but then i just have r\[^{3}\]

OpenStudy (turingtest):

1) multiply both sides by 3/4 2) divide both sides by \(\pi\) 3) take the cube root of both sides

OpenStudy (anonymous):

i forgot how to do cube root....I ended up with .75V/\[\pi\]=r cubed

OpenStudy (turingtest):

just write it out\[\frac3{4\pi}V=r^3\]\[r=\sqrt[3]{\frac3{4\pi}V}\]

OpenStudy (anonymous):

Ohhh now that makes sense now! :D you're really smart!